# Marginally interpretable spatial logistic regression with bridge processes

### v2

@@ A.2.3 Step 3 and additional remarks
When there are additional unknown parameters $\vartheta$, such as parameters of the correlation kernel, having prior $p(\vartheta)$, a sampling step for $\vartheta$ can be included in Step 2(i) with a suitable proposal $\vartheta^{\star}\sim q(\vartheta^{\star}\mid\vartheta)$. This is beneficial for mixing, since random effects $u_{1:n}$ are integrated out in Step 2(i). For the choice of proposal distribution in Step 2(i), we transform the parameter space to the Euclidean space and apply the adaptive scheme of Haario et al., (2001) based on a multivariate normal proposal.
@@ A.2.4 Details of scalable computation with low-rank dependence structure
We describe in detail how Algorithm 1 becomes scalable with the correlation kernel (8) with a low-rank structure. Denoting $R_{nq}=[\mathcal{R}(s_{i},\tilde{s}_{k})]_{i=1,k=1}^{n,q}$ as an $n\times q$ matrix and $R_{qn}=R_{nq}^{\mathrm{\scriptscriptstyle T}}$, we have a low-rank structured correlation matrix for $n$ realizations $\tilde{R}=[\tilde{R}(s_{i},s_{i^{\prime}})]_{i,i^{\prime}=1}^{n,n}=R_{nq}R_{qq}^{-1}R_{qn}+D_{nn}$ where $D_{nn}$ is a diagonal matrix with elements $1-r(s_{i})^{\mathrm{\scriptscriptstyle T}}R_{qq}^{-1}r(s_{i})$ for $i=1,\dots,n$.
First, step 1 of Algorithm 1 involves $n\times n$ matrix inversion in equation (A.1). This can be reduced to a $q\times q$ matrix inversion problem using the Woodbury matrix identity,
$(\lambda^{-1}\tilde{R}^{-1}+\Omega_{nn})^{-1}$ $=\lambda\{(R_{nq}R_{qq}^{-1}R_{qn}+D_{nn})^{-1}+\lambda\Omega_{nn}\}^{-1}$
$=\lambda\Delta_{1}+\lambda\Delta_{1}D_{nn}^{-1}R_{nq}(R_{qq}+R_{qn}\Delta_{2}R_{nq})^{-1}R_{qn}D_{nn}^{-1}\Delta_{1}$ (A.4)
where $\Delta_{1}=(\lambda\Omega_{nn}+D_{nn}^{-1})^{-1}$, $\Delta_{2}=D_{nn}^{-1}-D_{nn}^{-1}\Delta_{1}D_{nn}^{-1}$ are diagonal matrices.
Next, Step 2(i) involves $n$-dimensional normal density evaluation with covariance $\lambda\tilde{R}+\Omega_{nn}^{-1}$. Its inverse and determinant can be efficiently calculated as
$(\lambda\tilde{R}+\Omega_{nn}^{-1})^{-1}$ $=(\lambda R_{nq}R_{qq}^{-1}R_{qn}+\lambda D_{nn}+\Omega_{nn}^{-1})^{-1}$
$=\Delta_{3}-\Delta_{3}R_{nq}(\lambda^{-1}R_{qq}+R_{qn}\Delta_{3}R_{nq})^{-1}R_{qn}\Delta_{3},$
$|\lambda\tilde{R}+\Omega_{nn}^{-1}|$ $=|\Delta_{3}^{-1}|\times|\lambda R_{qq}^{-1}|\times|\lambda^{-1}R_{qq}+R_{qn}\Delta_{3}R_{nq}|$
where $\Delta_{3}=(\lambda D_{nn}+\Omega_{nn}^{-1})^{-1}$ is a diagonal matrix.
Finally, step 2(ii) corresponds to sampling $(u_{1:n}\mid\lambda,\omega,\beta)\sim\mathrm{N}_{n}(V_{2}b_{2},V_{2})$ with $V_{2}=(\Omega_{nn}+\lambda^{-1}\tilde{R}^{-1})^{-1}$ and $b_{2}=Z^{\mathrm{\scriptscriptstyle T}}\{(y-0.51_{N})-\Omega X\beta\}$. Using the expression (A.4) for $V_{2}$, $u_{1:n}$ can be sampled from the following steps: (1) Sample from $q$-dimensional multivariate normal $\tilde{u}\sim\mathrm{N}_{q}\{0,(R_{qq}+R_{qn}\Delta_{2}R_{nq})^{-1}\}$, (2) Sample $\epsilon\sim\mathrm{N}_{n}(0,\lambda\Delta_{1})$, (3) Set $u_{1:n}=V_{2}b_{2}+\lambda^{1/2}\Delta_{1}D_{nn}^{-1}R_{nq}\tilde{u}+\epsilon$, which reduces to sampling from a $q$-dimensional multivariate normal instead of an $n$-dimensional one.
@@ A.3 Details of multivariate bridge distribution
[…]
=$\stackrel{{\scriptstyle(**)}}{{=}}D(\phi)\sum_{k=1}^{\infty}(-1)^{k+1}C_{k}(\phi)\frac{\Gamma(d/2+1/2)2^{(d+1)/2}}{\{\pi^{2}C_{k}(\phi)^{2}/\phi^{2}+u^{\mathrm{\scriptscriptstyle T}}R^{-1}u\}^{(d+1)/2}}$
@@
where ()(*)( ∗ ) holds for $d\geq 2$ due to Fubini, and ()(**)( ∗ ∗ ) is from recognizing the integrand as an inverse gamma density with parameters $(d+1)/2$ and $\{\pi^{2}C_{k}(\phi)^{2}/\phi^{2}+u^{\mathrm{\scriptscriptstyle T}}R^{-1}u\}/2$. ∎
@@
A.4 Additional figure for Gambia malaria study

### v3

@@ A.2.3 Step 3 and additional remarks
When it comes to the empirical Bayes approach, where $\phi$ is fixed at $\hat{\phi}$, there are fewer benefits of using the particle marginal Metropolis-Hastings algorithm. In this case, we suggest sampling $\rho$ given $\lambda$ (Stage 2 - Step 2 of Algorithm 1) and $\lambda$ given $\rho$ (Stage 2 - Step 3 of Algorithm 1) using the Metropolis-Hastings algorithm based on expression (A.3), where we use an adaptive Metropolis-Hastings proposal for $\rho$ and Metropolis-Hastings for $\lambda$ with prior ${p_{\textsc{m}}}(\lambda;\phi)$ as an independent proposal distribution, which avoids evaluation of ${p_{\textsc{m}}}(\lambda;\phi)$ as they cancel out in the acceptance probability.
@@ A.2.4 Details of scalable computation with low-rank dependence structure
We describe in detail how Algorithm 1 becomes scalable with the correlation kernel (12) with a low-rank structure. Denoting $\mathbf{R}_{nq}=[\mathcal{R}(s_{i},\tilde{s}_{k})]_{i=1,k=1}^{n,q}$ as an $n\times q$ matrix and $\mathbf{R}_{qn}=\mathbf{R}_{nq}^{\mathrm{\scriptscriptstyle T}}$, we have a low-rank structured correlation matrix for $n$ realizations $\tilde{\mathbf{R}}=[\tilde{\mathcal{R}}(s_{i},s_{i^{\prime}})]_{i,i^{\prime}=1}^{n,n}=\mathbf{R}_{nq}\mathbf{R}_{qq}^{-1}\mathbf{R}_{qn}+\mathbf{D}_{nn}$ where $\mathbf{D}_{nn}$ is a diagonal matrix with elements $1-\mathbf{r}(s_{i})^{\mathrm{\scriptscriptstyle T}}\mathbf{R}_{qq}^{-1}\mathbf{r}(s_{i})$ for $i=1,\dots,n$.
First, step 1 of Algorithm 1 or Algorithm 2 involves $n\times n$ matrix inversion in equation (A.1). This can be reduced to a $q\times q$ matrix inversion problem using the Woodbury matrix identity,
$(\lambda^{-1}\tilde{\mathbf{R}}^{-1}+\bm{\Omega}_{nn})^{-1}$ $=\lambda\{(\mathbf{R}_{nq}\mathbf{R}_{qq}^{-1}\mathbf{R}_{qn}+\mathbf{D}_{nn})^{-1}+\lambda\bm{\Omega}_{nn}\}^{-1}$
$=\lambda\bm{\Delta}_{1}+\lambda\bm{\Delta}_{1}\mathbf{D}_{nn}^{-1}\mathbf{R}_{nq}(\mathbf{R}_{qq}+\mathbf{R}_{qn}\bm{\Delta}_{2}\mathbf{R}_{nq})^{-1}\mathbf{R}_{qn}\mathbf{D}_{nn}^{-1}\Delta_{1}$ (A.4)
where $\bm{\Delta}_{1}=(\lambda\bm{\Omega}_{nn}+\mathbf{D}_{nn}^{-1})^{-1}$, $\bm{\Delta}_{2}=\mathbf{D}_{nn}^{-1}-\mathbf{D}_{nn}^{-1}\bm{\Delta}_{1}\mathbf{D}_{nn}^{-1}$ are diagonal matrices.
Next, step 2, step 3 of Algorithm 1, and step 2(i) of Algorithm 2 involve $n$-dimensional normal density evaluation with covariance $\lambda\tilde{\mathbf{R}}+\bm{\Omega}_{nn}^{-1}$. Its inverse and determinant can be efficiently calculated as
$(\lambda\tilde{\mathbf{R}}+\bm{\Omega}_{nn}^{-1})^{-1}$ $=(\lambda\mathbf{R}_{nq}\mathbf{R}_{qq}^{-1}\mathbf{R}_{qn}+\lambda\mathbf{D}_{nn}+\bm{\Omega}_{nn}^{-1})^{-1}$
$=\bm{\Delta}_{3}-\bm{\Delta}_{3}\mathbf{R}_{nq}(\lambda^{-1}\mathbf{R}_{qq}+\mathbf{R}_{qn}\bm{\Delta}_{3}\mathbf{R}_{nq})^{-1}\mathbf{R}_{qn}\bm{\Delta}_{3},$
$|\lambda\tilde{\mathbf{R}}+\bm{\Omega}_{nn}^{-1}|$ $=|\bm{\Delta}_{3}^{-1}|\times|\lambda\mathbf{R}_{qq}^{-1}|\times|\lambda^{-1}\mathbf{R}_{qq}+\mathbf{R}_{qn}\Delta_{3}\mathbf{R}_{nq}|$
where $\bm{\Delta}_{3}=(\lambda\mathbf{D}_{nn}+\bm{\Omega}_{nn}^{-1})^{-1}$ is a diagonal matrix.
Finally, step 2(ii) corresponds to sampling $(\mathbf{u}_{1:n}\mid\lambda,\bm{\omega},\bm{\beta})\sim\mathrm{N}_{n}(\mathbf{V}_{2}\mathbf{b}_{2},\mathbf{V}_{2})$ with $\mathbf{V}_{2}=(\bm{\Omega}_{nn}+\lambda^{-1}\tilde{\mathbf{R}}^{-1})^{-1}$ and $\mathbf{b}_{2}=\mathbf{Z}^{\mathrm{\scriptscriptstyle T}}\{(\mathbf{y}-0.5\bm{1}_{N})-\bm{\Omega}\mathbf{X}\bm{\beta}\}$. Using the expression (A.4) for $\mathbf{V}_{2}$, $\mathbf{u}_{1:n}$ can be sampled from the following steps: (1) Sample from $q$-dimensional multivariate normal $\tilde{\mathbf{u}}\sim\mathrm{N}_{q}\{\bm{0},(\mathbf{R}_{qq}+\mathbf{R}_{qn}\Delta_{2}\mathbf{R}_{nq})^{-1}\}$, (2) Sample $\epsilon\sim\mathrm{N}_{n}(\bm{0},\lambda\bm{\Delta}_{1})$, (3) Set $\mathbf{u}_{1:n}=\mathbf{V}_{2}\mathbf{b}_{2}+\lambda^{1/2}\bm{\Delta}_{1}\mathbf{D}_{nn}^{-1}\mathbf{R}_{nq}\tilde{\mathbf{u}}+\epsilon$, which reduces to sampling from a $q$-dimensional multivariate normal instead of an $n$-dimensional one.
@@ A.3 Details of multivariate bridge distribution
$\stackrel{{\scriptstyle(*)}}{{=}}D(\phi)\sum_{k=1}^{\infty}(-1)^{k+1}C_{k}(\phi)\int_{0}^{\infty}\frac{1}{\lambda^{(d+3)/2}}\exp\left[-\frac{1}{2\lambda}\left\{\frac{\pi^{2}C_{k}(\phi)^{2}}{\phi^{2}}+u^{\mathrm{\scriptscriptstyle T}}R^{-1}u\right\}\right]d\lambda$
$\stackrel{{\scriptstyle(**)}}{{=}}D(\phi)\sum_{k=1}^{\infty}(-1)^{k+1}C_{k}(\phi)\frac{\Gamma(d/2+1/2)2^{(d+1)/2}}{\{\pi^{2}C_{k}(\phi)^{2}/\phi^{2}+u^{\mathrm{\scriptscriptstyle T}}R^{-1}u\}^{(d+1)/2}}$
@@
where ()(*) holds for $d\geq 2$ due to Fubini, and ()(**) is from recognizing the integrand as an inverse gamma density with parameters $(d+1)/2$ and $\{\pi^{2}C_{k}(\phi)^{2}/\phi^{2}+u^{\mathrm{\scriptscriptstyle T}}R^{-1}u\}/2$. ∎
@@
A.4 Additional results for simulation study
Table A.1: Comparison of estimated population-averaged effects based on 200 simulations in terms of bias, root mean squared error (RMSE), average length of 95% confidence or credible interval (CI.95), and coverage probabilities at nominal level 0.95 (Cover).
Population-averaged ($\hat{\beta}_{0}^{\textsc{m}}$) Data from bridge process Data from Gaussian copula
$\phi=0.7$ $\phi=0.9$ $\phi=0.7$ $\phi=0.9$
Setting Method Bias RMSE Bias RMSE Bias RMSE Bias RMSE
$\rho=0.05$ SpGEE 0.001 0.237 -0.019 0.165 -0.005 0.227 -0.024 0.147
BrP -0.003 0.231 -0.019 0.168 -0.003 0.235 -0.023 0.151
BrP-FB -0.003 0.245 -0.022 0.169 -0.005 0.237 -0.021 0.135
$\rho=0.1$ SpGEE -0.013 0.386 -0.018 0.298 -0.011 0.380 -0.034 0.231
BrP -0.001 0.028 -0.024 0.310 -0.002 0.404 -0.032 0.247
BrP-FB -0.002 0.382 -0.027 0.299 0.003 0.377 -0.025 0.220
SpGEE, spatial generalized estimating equation method of Cattelan and Varin, (2018); BrP, proposed bridge process model; BrP-FB, bridge process model with fully Bayesian approach on $\phi$.
Table A.2: Comparison of estimated site-specific effects based on 200 simulations in terms of bias, root mean squared error, average length of 95% confidence or credible interval, and coverage probabilities at nominal level 0.95.
Site-specific ($\hat{\beta}_{0}$) Data from bridge process Data from Gaussian copula
$\phi=0.7$ $\phi=0.9$ $\phi=0.7$ $\phi=0.9$
Setting Method Bias RMSE Bias RMSE Bias RMSE Bias RMSE
$\rho=0.05$ GP -0.022 0.412 -0.041 0.274 -0.005 0.338 -0.026 0.171
BrP -0.022 0.401 -0.040 0.278 -0.004 0.337 -0.025 0.168
BrP-FB -0.021 0.413 -0.040 0.276 -0.007 0.339 -0.026 0.170
$\rho=0.1$ GP -0.019 0.617 -0.057 0.505 0.006 0.568 -0.035 0.295
BrP -0.018 0.616 -0.056 0.497 0.000 0.560 -0.036 0.275
BrP-FB -0.019 0.621 -0.059 0.516 0.008 0.568 -0.033 0.293
SpGEE, spatial generalized estimating equation method of Cattelan and Varin, (2018); BrP, proposed bridge process model; BrP-FB, bridge process model with fully Bayesian approach on $\phi$.
A.5 Additional results for Gambia malaria data study
To better motivate the need for the spatial logistic model in analyzing the Gambia data, we analyzed the spatial dependence of the residuals through the variogram and Moran’s I test (Wikle et al., , 2019). We use village-level Pearson residual, defined as
$r_{i}=\frac{1}{\sqrt{N_{i}}}\sum_{j=1}^{N_{i}}\frac{Y_{ij}-\hat{p}_{ij}}{\sqrt{\hat{p}_{ij}(1-\hat{p}_{ij})}}$
for village $i$ ($\hat{p}_{ij}$ is an estimated probability of $Y_{ij}$), described in Section 3.1. of Diggle et al., (2002). Specifically, we analyze the village-level Pearson residuals under three models: (A) the logistic random intercept model (non-spatial), (B) the bridge process random effect logistic model, and (C) the Gaussian process random effect logistic model.
Figure A.4: Empirical variograms of village-level residuals $r_{i}$ for three logistic regression models (distance in km). (Left, A) Non-spatial random-intercept logistic model, which exhibits a clear increasing trend with distance. (Center, B) Bridge-process random-effect model, and (Right, C) Gaussian-process random-effect model, both of which show no evident spatial pattern. For (A), the fitted exponential variogram model is overlaid. For (B) and (C), fitting an exponential variogram model did not converge.
The empirical variograms based on residuals from the three models are displayed in Figure A.4. The variogram for model (A) shows a clear increasing pattern with distance, indicating substantial residual spatial dependence and motivating the need for spatial modeling. In contrast, the variograms for models (B) and (C) exhibit no discernible spatial structure, suggesting that the spatial random effects in these models successfully account for the underlying spatial dependence.
Moran’s I test using a 4-nearest-neighbor weight matrix further supports these observations. For model (A), the test yields $I=0.334$ ($p\text{-value}<10^{-5}$), providing strong evidence against the null hypothesis of no spatial correlation. For models (B) and (C), the results are I=0.193I=-0.193 ($p\text{-value}=0.987$) and I=0.202I=-0.202 ($p\text{-value}=0.991$), respectively, indicating no evidence of residual spatial dependence.
@@ A.5 Additional results for Gambia malaria data study
Finally, Figure A.5 shows MCMC trace plots of model parameters from 3 different chains, illustrating good mixing for both the bridge process random effect model and the Gaussian process random effect model.